OpenStudy (della):

mc015-1.jpg Find the height h of the parallelogram. 15 18 21 12

sam (.sam.):

Where is the diagram?

OpenStudy (mathmate):

Just a guess that the figure looks like this: 

OpenStudy (mathmate):

from the given information, we can deduce the following:  consequently prove that triangles AEB and DFA are similar. By proportionality of similar triangles, show that h/AB=AF/AD, or h/18=10/12 Cross multiply to find h.